OpenStudy (anonymous):

What is the exact value of tan(pi/6)?

OpenStudy (anonymous):

\[\tan \pi/6\]

OpenStudy (lgbasallote):

\[\Large \tan (\frac \pi 6) \implies \frac{\sin (\frac \pi 6)}{\cos (\frac \pi 6)}\] does that help?

OpenStudy (anonymous):

i think so. Wouldn't that turn into \[(\sqrt{3}/2)/(1/2)\]

OpenStudy (anonymous):

so the answer would be \[\sqrt{3}\]

hero (hero):

Well, nevermind, that's another way to do it

OpenStudy (lgbasallote):

hmm.. \[\sin(\frac \pi 6) = \frac 12\]isnt it?

hero (hero):

You don't have to convert to sine and cosine to solve it

OpenStudy (anonymous):



OpenStudy (lgbasallote):

well yeah...there is a right triangle method...

OpenStudy (anonymous):

you take the tangent of tan(pi/6)

OpenStudy (anonymous):

Whats the right triangle method?

OpenStudy (lgbasallote):

you remember your special triangles right? 30-60-90 triangle